What sum of two numbers is 67 and their difference is 25
Answer by oberobic(2304)

x+y = 67
so
x = 67-y
.
x-y = 25
so
x = 25+y
.
x=x
so
67-y = 25+y
67 = 25 + 2y
42 = 2y
2y = 42
y = 21
.
x = 67-21
x = 46
.
Check difference.
46-21=25
Correct.
.
Answer: The two numbers are 21 and 46.
.
Done.